I could really use some help on this problem, please try to help me solve this problem, I'm very desperate and I'm not great in math!
Mr. Huang operates a soybean farm outside of Grinnell, Iowa. To keep costs down, he buys many products in bulk and transfers them to smaller containers for use on the farm. Often the bulk products are not the correct concentration and need to be custom mixed before Mr. Huang can use them. One day he wants to apply fertilizer to a large field. A solution of 59% fertilizer is to be mixed with a solution of 38% fertilizer to form 42 liters of a 50% solution. How much of the 59% solution must he use?

Answer by mananth(16949)

percent ---------------- quantity
59 ---------------- x liters
38 ---------------- 42-x liters
50.00% ---------------- 42 liters
59x+38(42-x)=50*42
59x+1596 -38x=2100
59x-38x= 2100-1596
21x-=504
/21
x=24 liters 59.00%